Find the length of a side of a cube having a mass of 3.2 kg and the density of nuclear matter. 


1
Expert's answer
2021-06-29T13:07:31-0400

take the density of nuclear matter"=2.3\times 10^{17}kg/m^3"

"volume, V=\frac{mass}{density}=\frac{3.2}{2.3\times 10^{17}}=1.39\times 10^{-17}m^3"

"l=({1.39\times 10^{-17}})^{1/3}=2.405\times 10^{-6}m"